# Expressive Temporal Specifications for Reward Monitoring

### v2

@@ 2 Background
Reinforcement Learning (RL) environments are typically modelled as Markov decision processes (MDPs), defined as tuples $\mathcal{M}=(\mathcal{S},A,\Pr,\mathcal{R},\gamma)$, where $\mathcal{S}$ is a set of states; AA is a set of actions; $\Pr:\mathcal{S}\times A\to\Delta(\mathcal{S})$ is the transition kernel, with $\Delta(\mathcal{S})$ being the set of probability distributions over $\mathcal{S}$, $\mathcal{R}:\mathcal{S}\times A\times\mathcal{S}\to\mathbb{R}$ is the (Markovian) reward function, and $\gamma\in[0,1]$ is the discount factor.
@@
[…]
@@ 3.3 Monitor Synthesis
  • For temporal until $\varphi=\varphi_{1}\>\mathcal{U}\>\varphi_{2}$, let us start by obtaining $\mathcal{A}^{\varphi_{1}}$ and $\mathcal{A}^{\varphi_{2}}$ from procedure Synth. We form a new monitor $\mathcal{A}^{\varphi}$ as follows, comprising a state space of $Q=Q^{\varphi_{1}}\times Q^{\varphi_{2}}=\{(q^{\varphi_{1}},q^{\varphi_{2}})\mid q^{\varphi_{1}}\in Q^{\varphi_{1}},\,q^{\varphi_{2}}\in Q^{\varphi_{2}}\}$, and $q_{0}\;\leftarrow\;(q_{0}^{\varphi_{1}},\;q_{0}^{\varphi_{2}})$. Then, we can state $\mathcal{V}\leftarrow\mathcal{V}^{\varphi_{1}}\>\cup\>\mathcal{V}^{\varphi_{2}}\cup\bigl\{(t_{\min\varphi_{1}},1)\bigr\}\cup\bigl\{(t_{\max\varphi_{2}},0)\bigr\}\cup\bigl\{(t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}},0)\bigr\}$. Then, for all states $q\in Q$, we can assign $\delta_{v}(q,\mathcal{V}(t_{\min\varphi_{1}}))\leftarrow\min(\mathcal{V}(t_{\min\varphi_{1}}),\mathcal{V}(t_{\varphi_{1}}))$, $\delta_{v}(q,\mathcal{V}(t_{\max\varphi_{2}}))\leftarrow\max(\mathcal{V}(t_{\max\varphi_{2}}),\mathcal{V}(t_{\varphi_{2}}))$, and $\delta_{v}(q,t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}})\leftarrow\max(\mathcal{V}(t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}}),\min(\mathcal{V}(t_{\min\varphi_{1}}),\mathcal{V}(t_{\max\varphi_{2}})))$. Finally, we can assign $\delta_{r}\leftarrow\mathcal{V}(t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}})$.
@@
Monitors for temporal operators $F$ (Eventually) and $G$ (Always) can also be constructed primitively, with fewer registers as an implementation optimisation, only needing to track the max and min resp. of the associated subformula. Further details of optimisations and abbreviation connectives are provided in Appendix B.1.
@@
Let $\varphi$ be an ltlf[][$\mathcal{F}$] formula, $\mathcal{A}_{\varphi}$ the QRM constructed from $\varphi$ as per Alg. 1, and $\lambda$ a finite trace of length $n$. Let $\mathrm{val}_{\mathcal{A}_{\phi}}(\,\varphi,\,i\,)$ denote the value stored in the reward register at time index ii. Then, for each index $1\leq i\leq n$,
@@ 4 Experimental Evaluation
with span NN. Define checkpoints every NN steps: $C_{i}\stackrel{{\scriptstyle def}}{{=}}E_{iN}$.
@@
Figure 2: Task-completion percentages across safety-gridworlds and toy environments (also detailed in Table 1).
@@ 6 Conclusions and Future Work
In this paper, we have shown how to construct reward monitors using quantitative, linear time ltlf[][$\mathcal{F}$] specifications, and then demonstrated the effectiveness of our approach against Boolean monitors and handcrafted reward functions, across various environments with differing reward characteristics. We empirically showed that our quantitative monitor surpasses or matches the performance of a Boolean monitor, and in some cases, surpasses the performance of the manually-specified reward function. Learning uses the product between the environment MDP and the composite reward monitor formed from all specification-reward pairs. Additionally, we syntactically identify which monitors target safety properties, which when violated, have the power to block rewards received from all other monitors indefinitely. Code is available at https://github.com/nightly/quantitative-reward-monitoring.
@@ B.1 Construction
Since quantitative monitors have a deterministic transition function, we can let $q_{-1}$ and $q_{-2}$ denote the final state of a monitor and the penultimate state of a monitor respectively.
@@
    • If $\psi=F\varphi$, then $\mathcal{A}^{\varphi}\leftarrow\textsc{Synth}(\varphi)$, $Q\leftarrow Q^{\varphi}$, $q_{0}\leftarrow q_{0}^{\varphi}$, $\delta_{q}\leftarrow\delta_{q}^{\varphi}$, $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\varphi}\cup\{(t_{G(F\varphi)},0)\}$, and $\delta_{v}\leftarrow\delta_{v}^{\varphi}$. Then for all $q\in Q$, do $\delta_{v}(q,t_{G(F\varphi)})\leftarrow\mathcal{V}(t_{\varphi})$. And finally, $\delta_{r}\leftarrow\mathcal{V}(t_{G(F\varphi)})$.
    • Otherwise, $\mathcal{A}^{\psi}\leftarrow\textsc{synth}(\psi)$, $Q\leftarrow Q^{\psi}$, $q_{0}\leftarrow q_{0}^{\psi}$, $\delta_{q}\leftarrow\delta_{q}^{\psi}$, $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\psi}\cup\{(t_{G\psi},1)\}$, $\delta_{v}\leftarrow\delta_{v}^{\psi}$. Then, for all $q\in Q$ do: $\delta_{v}(q,t_{G\psi})\leftarrow\min(\mathcal{V}(t_{G_{\psi}}),\mathcal{V}(t_{\psi}))$. Finally, $\delta_{r}\leftarrow\mathcal{V}(t_{G\psi})$.
@@
    • For the case of $\psi=G\varphi$, then: $\mathcal{A}^{\varphi}\leftarrow\textsc{synth}(\varphi)$, $Q\leftarrow Q^{\varphi}$, $q_{0}\leftarrow q_{0}^{\varphi}$, $\delta_{q}\leftarrow\delta_{q}^{\varphi}$, $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\varphi}\cup\{(t_{F(G\varphi)},0)\}$, $\delta_{v}\leftarrow\delta_{v}^{\varphi}$. Then for all $q\in Q$, do $\delta_{v}(q,t_{F(G\varphi)})\leftarrow\mathcal{V}(t_{\varphi})$. Finally, $\delta_{r}\leftarrow\mathcal{V}(t_{F(G\psi)})$.
    • Otherwise, $\mathcal{A}^{\psi}\leftarrow\textsc{synth}(\psi)$, $Q\leftarrow Q^{\psi}$, $q_{0}\leftarrow q_{0}^{\psi}$, $\delta_{q}\leftarrow\delta_{q}^{\psi}$, $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\psi}\cup\{(t_{F\psi},1)\}$, $\delta_{v}\leftarrow\delta_{v}^{\psi}$. Then, for all $q\in Q$ do: $\delta_{v}(q,t_{F\psi})\leftarrow\max(\mathcal{V}(t_{F\psi}),\mathcal{V}(t_{\psi}))$. Finally, $\delta_{r}\leftarrow\mathcal{V}(t_{F\psi})$.
In the case where always $G$ and eventually $F$ are nested into each other, such that $F(G(\varphi))$ or $G(F(\varphi))$, this is simplified into $last(\varphi)$ by the ltlf[][$\mathcal{F}$] semantics. As we are using runtime monitoring, the current trace index is also technically the last index.
@@ B.4 Proof of Correctness
    $\varphi=p$, for an atomic proposition, the semantic value at a given time step $i$ is $\bigl[\![\,p,i\,]\!\bigr](\lambda)$. The monitor constructed transitions from the initial state to an empty self-looping state with no further register updates. The initial state provides a reward output of the current valuation of $p$ from the labelling function $\mathcal{L}_{p}(s)$ of the MDP, and this is stored permanently in the reward register, which corresponds to evaluating a proposition $p$ at index ii.
@@
    Conjunction $\varphi=\psi_{1}\land\psi_{2}$, by constructing the respective submonitors of $\mathcal{A}^{\psi}1$ and $\mathcal{A}^{\psi}2$, taking the Cartesian product of the state and transition space, tracking the propositional values of each formula, we append an additional register that tracks the minimum between the two formulae such that $\min(\psi_{1},\psi_{2})$ and use this as the reward register, which provides the correct output for time step ii.
@@
    Next $\varphi=X\psi$ is almost identical to the base case of an atomic proposition, but requires accessing the next timestep $i+1$ instead of ii. Whilst requiring a future time step’s value is typically an issue with runtime verification, when performing an action we have the triple ($s,a,s^{\prime}$), therefore the labelling of $s^{\prime}$ is considered for $i+1$.
@@ Appendix C Environment Specification Details
In all experiments, we used an exponentially moving average convergence test (see Definition 8) with span $N=32$ episodes, forming checkpoints every NN episodes and declaring convergence once the last $P=5$ checkpoint intervals showed less than a 1 relative improvement in the EMA of returns. The relative tolerance was automatically adapted to the recent noise in the checkpoint sequence, clipped to the range $\rho_{\min},\rho_{\max}]=[0.002,0.02]$, and we did not enforce a minimum EMA level or exclude negative returns.

### v3

@@ Abstract
Codehttps://github.com/nightly/quantitative-reward-monitoring
@@ 2 Background
Reinforcement Learning (RL) environments are typically modelled as Markov decision processes (MDPs), defined as tuples $\mathcal{M}=(\mathcal{S},A,\Pr,\mathcal{R},\gamma)$, where $\mathcal{S}$ is a set of states; $A$ is a set of actions; $\Pr:\mathcal{S}\times A\to\Delta(\mathcal{S})$ is the transition kernel, with $\Delta(\mathcal{S})$ being the set of probability distributions over $\mathcal{S}$, $\mathcal{R}:\mathcal{S}\times A\times\mathcal{S}\to\mathbb{R}$ is the (Markovian) reward function, and $\gamma\in[0,1]$ is the discount factor.
@@
012 Register updates ($\delta_{v}$) in State 2: $a\leftarrow\mathcal{L}_{a}(s^{\prime})$; $\lnot a\leftarrow 1-\mathcal{V}(a);\;$ $\min(\lnot a)\leftarrow\min\{\min(\lnot a),\lnot a\};\;$ $b\leftarrow\mathcal{L}_{b}(s^{\prime});\;$ $\max(b)\leftarrow\max\{F(b),b\}$ $\max(a\land\max(b))\leftarrow\max\{\max(a\land\max(b)),a\land\max(b)\}$ $\cdots$
@@ 3.3 Monitor Synthesis
  • For temporal until $\varphi=\varphi_{1}\>\mathcal{U}\>\varphi_{2}$, let us start by obtaining $\mathcal{A}^{\varphi_{1}}$ and $\mathcal{A}^{\varphi_{2}}$ from procedure Synth. Let $q_{-1}$ and $q_{-2}$ denote the final state of a monitor and the penultimate state of a monitor respectively. We form a new monitor $\mathcal{A}^{\varphi}$ as follows, comprising a state space of $Q=Q^{\varphi_{1}}\times Q^{\varphi_{2}}=\{(q^{\varphi_{1}},q^{\varphi_{2}})\mid q^{\varphi_{1}}\in Q^{\varphi_{1}},\,q^{\varphi_{2}}\in Q^{\varphi_{2}}\}$, and $q_{0}\;\leftarrow\;(q_{0}^{\varphi_{1}},\;q_{0}^{\varphi_{2}})$. Then, we can state $\mathcal{V}\leftarrow\mathcal{V}^{\varphi_{1}}\>\cup\>\mathcal{V}^{\varphi_{2}}\cup\bigl\{(t_{\min\varphi_{1}},1)\bigr\}\cup\bigl\{(t_{\max\varphi_{2}},0)\bigr\}\cup\bigl\{(t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}},0)\bigr\}$. Then, for all states $q\in Q$, we can assign $\delta_{v}(q,t_{\min\varphi_{1}})\leftarrow\min(\mathcal{V}(t_{\min\varphi_{1}}),\mathcal{V}(t_{\varphi_{1}}))$, $\delta_{v}(q,t_{\max\varphi_{2}})\leftarrow\max(\mathcal{V}(t_{\max\varphi_{2}}),\mathcal{V}(t_{\varphi_{2}}))$, and $\delta_{v}(q,t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}})\leftarrow\max(\mathcal{V}(t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}}),\min(\mathcal{V}(t_{\min\varphi_{1}}),\mathcal{V}(t_{\max\varphi_{2}})))$. To support continuous updates, we assign $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2},t)\leftarrow\delta_{v}(q_{-1},t)\>\forall t$, and all contained atomic registers are renamed (to avoid name clashing), and at initialisation are filled with the initial state labels $\mathcal{L}_{p}(\mathcal{S}_{0})$, updatable with $\delta_{v}(q_{0},t_{p})\to\mathcal{L}_{p}(s^{\prime})$. Finally, we can assign $\delta_{r}\leftarrow\mathcal{V}(t_{\varphi_{1}\>\mathcal{U}\>\varphi_{2}})$.
@@
Monitors for temporal operators $F$ (Eventually) and $G$ (Always) can also be constructed primitively, with fewer registers as an implementation optimisation, only needing to track the max and min resp. of the associated subformula. Further details of optimisations and abbreviated connectives are provided in Appendix B.1.
@@
Let $\varphi$ be an ltlf[][$\mathcal{F}$] formula, $\mathcal{A}_{\varphi}$ the QRM constructed from $\varphi$ as per Alg. 1, and $\lambda$ a finite trace of length $n$. Let $\mathrm{val}_{\mathcal{A}_{\phi}}(\,\varphi,\,i\,)$ denote the value stored in the reward register at time index $i$. Then, for each index $1\leq i\leq n$,
@@ 4 Experimental Evaluation
with span $N$. Define checkpoints every $N$ steps: $C_{i}\stackrel{{\scriptstyle def}}{{=}}E_{iN}$.
@@
Figure 2: Task-completion percentages across safety-gridworlds and toy environments, $\pm$ 1 std. Also detailed in Table 1.
@@ 6 Conclusions and Future Work
In this paper, we have shown how to construct reward monitors using quantitative, linear time ltlf[][$\mathcal{F}$] specifications, and then demonstrated the effectiveness of our approach against Boolean monitors and handcrafted reward functions, across various environments with differing reward characteristics. We empirically showed that our quantitative monitor surpasses or matches the performance of a Boolean monitor, and in some cases, surpasses the performance of the manually-specified reward function. Learning uses the product between the environment MDP and the composite reward monitor formed from all specification-reward pairs. Additionally, we syntactically identify which monitors target safety properties, which when violated, have the power to block rewards received from all other monitors indefinitely.
@@ B.1 Construction
Since quantitative monitors have a deterministic transition function, we can let $q_{-1}$ and $q_{-2}$ denote the final state of a monitor and the penultimate state of a monitor respectively. As with temporal until $\mathcal{U}$, all contained atomic registers are initialised using the initial state labels $\mathcal{L}_{p}(\mathcal{S}_{0})$ and are updated on each step using the successor-state labels $\mathcal{L}_{p}(s^{\prime})$, and registers are renamed to avoid name clashing.
@@
    • If $\psi=F\varphi$, then $\mathcal{A}^{\varphi}\leftarrow\textsc{Synth}(\varphi)$, $Q\leftarrow Q^{\varphi}$, $q_{0}\leftarrow q_{0}^{\varphi}$, continuous updates are supported by using the renamed registers and adding a loop between the penultimate and ultimate state: $\delta_{q}\leftarrow\delta_{q}^{\varphi}$, $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\varphi}\cup\{(t_{G(F\varphi)},0)\}$, and $\delta_{v}\leftarrow\delta_{v}^{\varphi}$. Then for all $q\in Q$, do $\delta_{v}(q,t_{G(F\varphi)})\leftarrow\mathcal{V}(t_{\varphi})$. And finally, $\delta_{r}\leftarrow\mathcal{V}(t_{G(F\varphi)})$.
    • Otherwise, $\mathcal{A}^{\psi}\leftarrow\textsc{synth}(\psi)$, $Q\leftarrow Q^{\psi}$, $q_{0}\leftarrow q_{0}^{\psi}$, $\delta_{q}\leftarrow\delta_{q}^{\psi}$, continuous updates are supported by using the renamed registers and adding a loop between the penultimate and ultimate state: $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\psi}\cup\{(t_{G\psi},1)\}$, $\delta_{v}\leftarrow\delta_{v}^{\psi}$. Then, for all $q\in Q$ do: $\delta_{v}(q,t_{G\psi})\leftarrow\min(\mathcal{V}(t_{G_{\psi}}),\mathcal{V}(t_{\psi}))$. Finally, $\delta_{r}\leftarrow\mathcal{V}(t_{G\psi})$.
@@
    • For the case of $\psi=G\varphi$, then: $\mathcal{A}^{\varphi}\leftarrow\textsc{synth}(\varphi)$, $Q\leftarrow Q^{\varphi}$, $q_{0}\leftarrow q_{0}^{\varphi}$, $\delta_{q}\leftarrow\delta_{q}^{\varphi}$, continuous updates are supported by using the renamed registers and adding a loop between the penultimate and ultimate state: $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\varphi}\cup\{(t_{F(G\varphi)},0)\}$, $\delta_{v}\leftarrow\delta_{v}^{\varphi}$. Then for all $q\in Q$, do $\delta_{v}(q,t_{F(G\varphi)})\leftarrow\mathcal{V}(t_{\varphi})$. Finally, $\delta_{r}\leftarrow\mathcal{V}(t_{F(G\psi)})$.
    • Otherwise, $\mathcal{A}^{\psi}\leftarrow\textsc{synth}(\psi)$, $Q\leftarrow Q^{\psi}$, $q_{0}\leftarrow q_{0}^{\psi}$, $\delta_{q}\leftarrow\delta_{q}^{\psi}$, continuous updates are supported by using the renamed registers and adding a loop between the penultimate and ultimate state: $\delta_{q}(q_{-1})\leftarrow q_{-2}$, $\delta_{v}(q_{-2})\leftarrow\delta_{v}(q_{-1})$, $\mathcal{V}\leftarrow\mathcal{V}^{\psi}\cup\{(t_{F\psi},0)\}$, $\delta_{v}\leftarrow\delta_{v}^{\psi}$. Then, for all $q\in Q$ do: $\delta_{v}(q,t_{F\psi})\leftarrow\max(\mathcal{V}(t_{F\psi}),\mathcal{V}(t_{\psi}))$. Finally, $\delta_{r}\leftarrow\mathcal{V}(t_{F\psi})$.
In the case where always $G$ and eventually $F$ are nested into each other, such that $F(G(\varphi))$ or $G(F(\varphi))$, this is simplified into $last(\varphi)$ by the ltlf[][$\mathcal{F}$] semantics. As we are using runtime monitoring, the current trace index, $s^{\prime}$ from a transition tuple ($s,a,s^{\prime}$) corresponds to the last trace index.
@@ B.4 Proof of Correctness
    $\varphi=p$, for an atomic proposition, the semantic value at a given time step $i$ is $\bigl[\![\,p,i\,]\!\bigr](\lambda)$. The monitor constructed transitions from the initial state to an empty self-looping state with no further register updates. The initial state provides a reward output of the current valuation of $p$ from the labelling function $\mathcal{L}_{p}(s)$ of the MDP, and this is stored permanently in the reward register, which corresponds to evaluating a proposition $p$ at index $i$.
@@
    Conjunction $\varphi=\psi_{1}\land\psi_{2}$, by constructing the respective submonitors of $\mathcal{A}^{\psi}1$ and $\mathcal{A}^{\psi}2$, taking the Cartesian product of the state and transition space, tracking the propositional values of each formula, we append an additional register that tracks the minimum between the two formulae such that $\min(\psi_{1},\psi_{2})$ and use this as the reward register, which provides the correct output for time step $i$.
@@
    Next $\varphi=X\psi$ is almost identical to the base case of an atomic proposition, but requires accessing the next timestep $i+1$ instead of $i$. Whilst requiring a future time step’s value is typically an issue with runtime verification, when performing an action we have the triple ($s,a,s^{\prime}$), therefore the labelling of $s^{\prime}$ is considered for $i+1$.
@@ Appendix C Environment Specification Details
In all experiments, we used an exponentially moving average convergence test (see Definition 8) with span $N=32$ episodes, forming checkpoints every $N$ episodes and declaring convergence once the last $P=5$ checkpoint intervals showed less than a 1 relative improvement in the EMA of returns. The relative tolerance was automatically adapted to the recent noise in the checkpoint sequence, clipped to the range $\rho_{\min},\rho_{\max}]=[0.002,0.02]$, and we did not enforce a minimum EMA level or exclude negative returns.